# Decentralized Stochastic Subgradient-based Methods for Nonsmooth Nonconvex Optimization

### v2

@@ 5.2 Performance comparison between DSGD, DSGD-M, and DSignSGD
We present the numerical results of the above-mentioned decentralized stochastic subgradient-based methods in Figure 1 and 2, including the test accuracy, train loss, and consensus error. It is worth mentioning that the consensus error at the $k$-th iteration is measured by $\frac{1}{\sqrt{d}}\|{\bm{X}}_{k}({\bm{I}}_{d}-{\bm{P}})\|_{\mathrm{F}}$. As illustrated in Figure 1 and 2, all the compared decentralized stochastic subgradient-based methods are able to train the ResNet-50 network to a high accuracy. Moreover, we can conclude from Figure 1(a) and 2(a) that DSignSGD achieves similar test accuracy as DSGD and better test accuracy compared to DSGD-M, while showing superior performance to DSGD and DSGD-M in the aspect of train loss from Figure 1(b) and 2(b). Furthermore, in all the compared methods, the consensus errors diminish to 00 throughout the iterations. These results illustrate that our framework is able to yield efficient stochastic subgradient-based methods for decentralized training tasks.
@@ 6 Conclusion
To address these issues, we propose a unified framework in (DSM) for developing stochastic subgradient-based methods in nonsmooth decentralized optimization. We conduct consensus analysis and relate the iterates of (DSM) to the trajectories of its corresponding noiseless differential inclusion. Then, under mild conditions, we establish the global convergence of (DSM), in the sense that for any $i,j\in[d]$, $\lim_{k\to+\infty}\left\|z_{i,k}=z_{j,k}\right\|=0$ and $\{z_{i,k}\}$ converges to the stable set of the corresponding noiseless differential inclusion.
Moreover, we show our proposed framework (DSM) is general enough to enclose the implementations of various acceleration techniques for decentralized stochastic subgradient methods, including momentum and gradient-tracking techniques. In particular, we demonstrate that a wide range of decentralized stochastic subgradient-based methods, including DSGD, DSGD-M, and DSGD-T, fits into our proposed framework. Therefore, our theoretical results, for the first time, provide the global convergence guarantees for these decentralized stochastic subgradient-based methods in nonsmooth nonconvex optimization. Furthermore, based on our proposed framework, we develop a new method named DSignSGD, where sign-mapping is employed to regularize the update directions. Preliminary numerical results verify the validity of our developed theoretical results and exhibit the high efficiency of these decentralized stochastic subgradient-based methods enclosed in our framework (DSM).
Our work sets several directions for future research in the field of nonsmooth nonconvex decentralized optimization, including the convergence of decentralized methods with time-varying networks and decentralized methods that incorporate asynchronous communication or compressed communication. Additionally, we are also eager to explore routes that presume relaxed conditions for convergence analysis of decentralized methods in solving (DOP). For instance, closed-measure approaches [51, 52, 35] are developed for establishing the convergence properties of stochastic subgradient-based methods, with relaxed conditions on the step-sizes and evaluation noises. How to employ these closed-measure approaches for analyzing decentralized stochastic subgradient-based methods is worthy of investigating.
@@ Acknowledgement
The work of Professor Xin Liu was supported in part by the National Key R&D Program of China (2023YFA1009300), the National Natural Science Foundation of China (12125108, 12226008, 11991021, 11991020, 12021001, 12288201), Key Research Program of Frontier Sciences, Chinese Academy of Sciences (ZDBS-LY-7022), and CAS-Croucher Funding Scheme for Joint Laboratories “CAS AMSS-PolyU Joint Laboratory of Applied Mathematics: Nonlinear Optimization Theory, Algorithms and Applications”.

### v3

@@ 5.2 Performance comparison between DSGD, DSGD-M, and DSignSGD
We present the numerical results of the above-mentioned decentralized stochastic sub-gradient-based methods in Figure 1 and 2, including the test accuracy, train loss, and consensus error. It is worth mentioning that the consensus error at the $k$-th iteration is measured by $\frac{1}{\sqrt{d}}\|{\bm{X}}_{k}({\bm{I}}_{d}-{\bm{P}})\|_{\mathrm{F}}$. As illustrated in Figure 1 and 2, all the compared decentralized stochastic subgradient-based methods are able to train the ResNet-50 network to a high accuracy. Moreover, we can conclude from Figure 1(a) and 2(a) that DSignSGD achieves similar test accuracy as DSGD and better test accuracy compared to DSGD-M, while showing superior performance to DSGD and DSGD-M in the aspect of train loss from Figure 1(b) and 2(b). Furthermore, in all the compared methods, the consensus error diminishes to 00 throughout the iterations. These results illustrate that our framework is able to yield efficient stochastic subgradient-based methods for decentralized training tasks.
@@ 6 Conclusion
To address these issues, we propose a unified framework (DSM) for analyzing the global convergence of stochastic subgradient-based methods in nonsmooth decentralized optimization. We conduct consensus analysis and relate the average iterates of (DSM) to the trajectories of its corresponding noiseless differential inclusion. Then, under two different evaluation noise settings, we establish convergence properties of (DSM), in the sense that $\{{\bm{z}}_{i,k}\}$ asymptotically reaches consensus and converges to the stable set of the corresponding noiseless differential inclusion.
Moreover, we show our proposed framework (DSM) is general enough to enclose the implementations of various acceleration techniques for decentralized stochastic subgradient methods, including momentum and gradient-tracking techniques. In particular, we demonstrate that a wide range of decentralized stochastic subgradient-based methods, such as DSGD, DSGD-M, and DSGD-T, together with two common sampling schemes, are all encompassed by our proposed framework. Therefore, our theoretical results, for the first time, provide the global convergence guarantees for these decentralized stochastic subgradient-based methods in nonsmooth nonconvex optimization. Furthermore, based on our proposed framework, we develop a new method named DSignSGD, where sign-mapping is employed to regularize the update directions. Preliminary numerical results verify the validity of our developed theoretical results and exhibit the high efficiency of these decentralized stochastic subgradient-based methods enclosed in our framework (DSM).
Our work sets several directions for future research in the field of nonsmooth nonconvex decentralized optimization. For example, non-asymptotic convergence analyses have been established in several existing works [28, 30], but their proof techniques cannot be extended to non-Clarke-regular cases. How to establish the non-asymptotic convergence rate for (DSM) is worthy of investigating. Further attention is also needed on the convergence of decentralized methods with time-varying networks, as well as those that incorporate asynchronous or compressed communication. Additionally, we are also eager to explore routes that presume relaxed conditions for convergence analysis of decentralized methods in solving (DOP). For instance, closed-measure approaches [60, 44, 42] are developed for establishing the convergence properties of stochastic subgradient-based methods, with relaxed conditions on the step-sizes and evaluation noises. How to employ these closed-measure approaches for analyzing decentralized stochastic subgradient-based methods is worthy of investigating.
@@ Acknowledgement
The work of Professor Xin Liu was supported in part by the National Key R&D Program of China (2023YFA1009300), the National Natural Science Foundation of China (12125108, 12226008, 12021001, 12288201), Key Research Program of Frontier Sciences, Chinese Academy of Sciences (ZDBS-LY-7022), and CAS-Croucher Funding Scheme for Joint Laboratories “CAS AMSS-PolyU Joint Laboratory of Applied Mathematics: Nonlinear Optimization Theory, Algorithms and Applications”.
Appendix A Proofs of Theorem 4.5 and 4.8
In this part, we devote to showing the avoidance of spurious critical points for Algorithm 1 and 2. In other words, these algorithms can avoid points that are critical for a specific conservative field but are not Clarke-critical points. We begin with the definition of the concept of almost everywhere $\mathcal{C}^{1}$ set-valued mapping.
Definition A.1.
A measurable mapping $q:\mathbb{R}^{nd}\to\mathbb{R}^{nd}$ is almost everywhere $\mathcal{C}^{1}$ if for almost every ${\bm{z}}\in\mathbb{R}^{nd}$, $q$ is locally continuously differentiable in a neighborhood of $z$.
Moreover, a set-valued mapping $\mathcal{Q}:\mathbb{R}^{nd}\rightrightarrows\mathbb{R}^{nd}$ is almost everywhere $\mathcal{C}^{1}$ if there exists an almost everywhere $\mathcal{C}^{1}$ mapping $q:\mathbb{R}^{nd}\to\mathbb{R}^{nd}$ such that for almost every ${\bm{z}}\in\mathbb{R}^{nd}$, $\mathcal{Q}({\bm{z}})=\{q({\bm{z}})\}$.
A.1 Proof of Theorem 4.5
For ease of presentation, we define
$F_{i,\mathcal{B}_{i,k+1}}({\bm{x}}):=\frac{1}{|\mathcal{B}_{i,k+1}|}\sum_{{\bm{s}}_{l}\in\mathcal{B}_{i,k+1}}F_{i}({\bm{x}};{\bm{s}}_{l}),\quad\text{ and }\quad D_{F_{i},\mathcal{B}_{i,k+1}}({\bm{x}}):=\frac{1}{|\mathcal{B}_{i,k+1}|}\sum_{{\bm{s}}_{l}\in\mathcal{B}_{i,k+1}}D_{F_{i}}({\bm{x}};{\bm{s}}_{l}).$
Define the set-valued mappings $\mathcal{U}_{k}:\mathbb{R}^{nd}\rightrightarrows\mathbb{R}^{nd}$ as
$\mathcal{U}_{k}({\bm{x}})$ $:=\left(D_{F_{1},\mathcal{B}_{1,k+1}}({\bm{x}}_{(1)}),\ldots,D_{F_{d},\mathcal{B}_{d,k+1}}({\bm{x}}_{(d)})\right)^{\top}$
Note that $(F_{i}(\cdot,{\bm{s}}_{l}),D_{F_{i}}(\cdot,{\bm{s}}_{l}))$ has a $\mathcal{C}^{r}$ variational stratification [36, Theorem 4], hence the mappings $\{\mathcal{U}_{k}\}$ are almost everywhere $\mathcal{C}^{1}$ over $\mathbb{R}^{nd}$, denoted by $\mathcal{U}_{k}({\bm{x}})=\{u_{k}({\bm{x}})\},$ for a.e. ${\bm{x}}\in\mathbb{R}^{nd}$. Furthermore, let
$\mathcal{Y}$ :=$:=\left\{{\bm{x}}\in\mathbb{R}^{nd}:D_{F_{i}}({\bm{x}};{\bm{s}}_{l})\neq\{\nabla F_{i}({\bm{x}},{\bm{s}}_{l})\},\text{ for some }i\in[N]\text{ and some }{\bm{s}}_{l}\in\mathcal{S}_{i}\right\},$
which is a zero-measure subset.
With a slight abuse of notation, we vectorize the sequence $\{{\bm{X}}_{k}\}$ by column-major order, and denote as ${\bm{x}}_{k}\in\mathbb{R}^{nd}$. Then the vectorized sequence $\{{\bm{x}}_{k}\}$ generated by Algorithm 1 satisfies
${\bm{x}}_{k+1}$ $\in{}({\bm{W}}\otimes{\bm{I}}_{n}){\bm{x}}_{k}-c\hat{\eta}_{k}\mathcal{U}_{k}({\bm{x}}_{k}).$ (A.1)
Set ${\bm{P}}:={\bm{W}}\otimes{\bm{I}}_{n}$, $\mathcal{Q}_{k}:=\hat{\eta}_{k}\mathcal{U}_{k}$, and $q_{k}:=\hat{\eta}_{k}{u}_{k}$. The Jacobian of the mapping ${\bm{P}}{\bm{x}}-cq_{k}({\bm{x}})$ is
${\bm{J}}_{c,k}({\bm{x}})={\bm{W}}\otimes{\bm{I}}_{n}-c\hat{\eta}_{k}\mathrm{Diag}(\nabla^{2}F_{1,\mathcal{B}_{1,k+1}}({\bm{x}}_{(1)}),\ldots,\nabla^{2}F_{d,\mathcal{B}_{d,k+1}}({\bm{x}}_{(d)})).$ (A.2)
The determinant of ${\bm{J}}_{c,k}({\bm{x}})$ is a zero polynomial of $c$ if and only if there exists ${\bm{v}}=[{\bm{v}}_{1}^{\top},\ldots,{\bm{v}}_{d}^{\top}]^{\top}$ such that ${\bm{v}}\in\mathrm{Ker}({\bm{W}}\otimes{\bm{I}}_{n})$ and ${\bm{v}}_{i}\in\mathrm{Ker}(\nabla^{2}F_{i,\mathcal{B}_{i,k+1}}({\bm{x}}_{(i)}))$ for any $i\in[d]$. As each $F_{i}(\cdot,{\bm{s}}_{l})$ is a definable function, we can infer that such $({\bm{x}}_{(1)},\ldots,{\bm{x}}_{(d)})$ constitutes a zero-measure subset of $\mathbb{R}^{nd}$. We refer to $\Gamma_{k}$ as the full-measure subset of $\mathbb{R}^{nd}$, in which $\mathcal{U}_{k}({\bm{x}})=\{u_{k}({\bm{x}})\},$ and $\det({\bm{J}}_{c,k}({\bm{x}}))$ is a non-trival $nd$-th order polynomial of $c$. This indicates the set
$\{c\in\mathbb{R}:\det({\bm{J}}_{c,k}({\bm{x}}))=0\}$
is zero-measure in $\mathbb{R}$. According to Fubini’s Theorem, there exists a full-measure subset $\mathcal{S}_{k}\subseteq\mathbb{R}_{+}$, such that for any $c\in\mathcal{S}_{k}$, $\{{\bm{x}}\in\Gamma_{k}:\det({\bm{J}}_{c,k}({\bm{x}}))\neq 0\}$ is a full-measure subset in $\mathbb{R}^{nd}$.
Applying the inverse function theorem, we can conclude that for any $c\in\mathcal{S}_{k}$ and any ${\bm{x}}\in\{x\in\Gamma_{k}:\det({\bm{J}}_{c,k}({\bm{x}}))\neq 0\}$, ${\bm{P}}{\bm{x}}-cq_{k}({\bm{x}})$ is a local diffeomorphism in a neighborhood of ${\bm{x}}$, denoted by $\mathbb{B}({\bm{x}},\delta_{{\bm{x}}})$. As a result,
$\{({\bm{P}}-cq_{k})(\mathbb{B}({\bm{x}}^{i},\delta_{{\bm{x}}^{i}}))\}_{{\bm{x}}\in\{x\in\Gamma_{k}:\det({\bm{J}}_{c,k}({\bm{x}}))\neq 0\}}$
is an open covering of $\{{\bm{x}}\in\Gamma_{k}:\det({\bm{J}}_{c,k}({\bm{x}}))\neq 0\}$. Based on Lindelof’s lemma [69], there exists countable $\{{\bm{x}}^{i}\}_{i\in\mathbb{N}_{+}}\subseteq\{{\bm{x}}\in\Gamma_{k}:\det({\bm{J}}_{c,k}({\bm{x}}))\neq 0\}$, such that
$\{{\bm{x}}\in\Gamma_{k}:\det({\bm{J}}_{c,k}({\bm{x}}))\neq 0\}\subseteq\bigcup_{i\in\mathbb{N}_{+}}({\bm{P}}-cq_{k})(\mathbb{B}({\bm{x}}^{i},\delta_{{\bm{x}}^{i}})).$
Given any zero-measure subset $A\subseteq\mathbb{R}^{nd}$, since ${\bm{P}}-cq_{k}$ is a diffeomorphism over $\mathbb{B}({\bm{x}}^{i},\delta_{{\bm{x}}^{i}})$, it can be deduced that the set $\{{\bm{z}}\in\mathbb{B}({\bm{x}}^{i},\delta_{{\bm{x}}^{i}}),{\bm{P}}{\bm{z}}-cq_{k}({\bm{z}})\in A\}$ is zero-measure. Therefore, the set $\bigcup_{i\in\mathbb{N}_{+}}\{{\bm{z}}\in\mathbb{B}({\bm{x}}^{i},\delta_{{\bm{x}}^{i}}),{\bm{P}}{\bm{z}}-cq_{k}({\bm{z}})\in A\}$ is a zero-measure subset, which yields that
$\{{\bm{z}}\in\mathbb{R}^{nd}:{\bm{P}}{\bm{z}}-cq_{k}({\bm{z}})\in A\}\subset\bigcup_{i\in\mathbb{N}_{+}}\{{\bm{z}}\in\mathbb{B}({\bm{x}}^{i},\delta_{{\bm{x}}^{i}}),{\bm{P}}{\bm{z}}-cq_{k}({\bm{z}})\in A\}\cup\{{\bm{z}}\in\Gamma_{k}:\det({\bm{J}}_{c,k}({\bm{z}}))\neq 0\}^{c}$
is zero-measure. That is to say, for any $c\in\mathcal{S}_{k}$, the set $\{{\bm{z}}\in\mathbb{R}^{nd}:{\bm{P}}{\bm{z}}-cq_{k}({\bm{z}})\in A\}$ is zero-measure.
Let $\mathcal{S}:=\cap_{k\geq p}\mathcal{S}_{k}$, $A_{0}:=A=\mathcal{Y},A_{k+1}:=\{{\bm{x}}\in\mathbb{R}^{nd}:{\bm{P}}{\bm{z}}-cq_{k}({\bm{z}})\in A\}$, and $\hat{A}:=\cup_{k\geq 0}A_{k}$. It can be inferred that $\mathcal{S}$ and $\hat{A}^{c}$ are full-measure subsets in $\mathbb{R}$ and $\mathbb{R}^{nd}$, respectively. For any $c\in\mathcal{S}$, and any ${\bm{x}}\in\hat{A}^{c}$, we have ${\bm{P}}{\bm{x}}-cq_{k}({\bm{x}})\in A^{c}$. Employing Fubini’s theorem, the set $\{c\in\mathbb{R},{\bm{x}}\in\mathbb{R}^{nd},\cup_{k\geq 0}\{{\bm{P}}{\bm{x}}-cq_{k}({\bm{x}})\}\subseteq A^{c}\}$ is full-measure in ×$\mathbb{R}\times\mathbb{R}^{nd}$, and thus, there exists a full-measure subset $\hat{\mathcal{S}}\times\hat{\mathcal{K}}$, whenever $(c,{\bm{x}}_{0})\in\hat{\mathcal{S}}\times\hat{\mathcal{K}}$, it follows that $\cup_{k\geq 0}\{{\bm{P}}{\bm{x}}-cq_{k}({\bm{x}})\}\subseteq A^{c}$, i.e. $\{{\bm{x}}_{k}\}\in\mathcal{Y}^{c}$.
Furthermore, the update scheme of vectorized sequence $\{{\bm{x}}_{k}\}$ can be reformulated as
[…]
Therefore, by applying Theorem 4.4 with $D_{f}=\partial f$, we can choose $\mathcal{S}=\hat{\mathcal{S}}\cap(0,\alpha_{c})$ and $\mathcal{K}=\hat{\mathcal{K}}\cap{\mathcal{X}}_{0}$, such that the claim in Theorem 4.5 holds.
A.2 Proof of Theorem 4.8
Let the set-valued mappings $\mathcal{U}_{1,k}:\mathbb{R}^{2nd}\rightrightarrows\mathbb{R}^{nd}$, $\mathcal{U}_{2,k}:\mathbb{R}^{2nd}\rightrightarrows\mathbb{R}^{2nd}$ be defined as
$\mathcal{U}_{1,k}({\bm{x}},{\bm{y}})$ $:=\left(\partial\hat{\phi}({\bm{y}}),{\bm{0}}\right)$
$\mathcal{U}_{2,k}({\bm{x}},{\bm{y}})$ $:=\left({\bm{0}},\tau({\bm{W}}\otimes{\bm{I}}_{n}){\bm{y}}-\tau\left(D_{F_{1},\mathcal{B}_{1,k+1}}({\bm{x}}_{(1)}),\ldots,D_{F_{d},\mathcal{B}_{d,k+1}}({\bm{x}}_{(d)})\right)^{\top}\right)$
where $\hat{\phi}({\bm{y}}):=(\phi({\bm{y}}_{(1)}),\ldots,\phi({\bm{y}}_{(d)}))^{\top}$, $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is convex and admits a unique minimizer at ${\bm{0}}$. Notice that $(F_{i}(\cdot,{\bm{s}}_{l}),D_{F_{i}(\cdot,{\bm{s}}_{l})})$ has a $\mathcal{C}^{r}$ variational stratification [36, Theorem 4], and $\hat{\phi}$ is differentiable on a full-measure subset of $\mathbb{R}^{nd}$, thus the mappings $\{\mathcal{U}_{j,k}\},j=1,2$ are almost everywhere $\mathcal{C}^{1}$ over $\mathbb{R}^{2nd}$, the corresponding almost-everywhere-$\mathcal{C}^{1}$ mappings are denoted by $u_{j,k}({\bm{z}}),j=1,2$. Furthermore, let
$\mathcal{Y}$ :=$:=\left\{({\bm{x}},{\bm{y}})\in\mathbb{R}^{2nd}:\hat{\phi}({\bm{y}})\text{ is not differentiable },D_{F_{i,l}}({\bm{x}})\neq\{\nabla F_{i}({\bm{x}},{\bm{s}}_{l})\},\text{ for some }i\in[N]\right.$
[…]
which is a zero-measure subset in $\mathbb{R}^{2nd}$.
With the same notations in Section A.1, the vectorized sequence $\{({\bm{x}}_{k},{\bm{y}}_{k})\}$ generated by Algorithm 2 satisfies
$({\bm{x}}_{k+\frac{1}{2}},{\bm{y}}_{k+\frac{1}{2}})$ $\in{}({\bm{W}}\otimes{\bm{I}}_{n}\times{\bm{I}}_{nd})({\bm{x}}_{k},{\bm{y}}_{k})-c\hat{\eta}_{k}\mathcal{U}_{1,k}({\bm{x}}_{k},{\bm{y}}_{k}),$ (A.3)
$\left({\bm{x}}_{k+1},{\bm{y}}_{k+1}\right)$ $\in{}({\bm{I}}_{nd}\times{\bm{W}}\otimes{\bm{I}}_{n})({\bm{x}}_{k+\frac{1}{2}},{\bm{y}}_{k+\frac{1}{2}})-c\hat{\eta}_{k}\mathcal{U}_{2,k}({\bm{x}}_{k+\frac{1}{2}},{\bm{y}}_{k+\frac{1}{2}}),$
Set ${\bm{P}}_{1}:={\bm{W}}\otimes{\bm{I}}_{n}\times{\bm{I}}_{nd}$, ${\bm{P}}_{2}:={\bm{I}}_{nd}\times{\bm{W}}\otimes{\bm{I}}_{n}$, then the update scheme of vectorized sequence $\{({\bm{x}}_{k},{\bm{y}}_{k})\}$ is
$({\bm{x}}_{k+1},{\bm{y}}_{k+1})$ ${}={\bm{P}}_{2}({\bm{P}}_{1}({\bm{x}}_{k},{\bm{y}}_{k}))-c{\bm{P}}_{2}(\hat{\eta}_{k}\mathcal{U}_{1,k}({\bm{x}}_{k},{\bm{y}}_{k}))-c\hat{\eta}_{k}\mathcal{U}_{2,k}({\bm{P}}_{1}({\bm{x}}_{k},{\bm{y}}_{k})-c\hat{\eta}_{k}\mathcal{U}_{1,k}({\bm{x}}_{k},{\bm{y}}_{k}))$
${}:={\bm{P}}({\bm{x}}_{k},{\bm{y}}_{k})-c\mathcal{Q}_{k}({\bm{x}}_{k},{\bm{y}}_{k})$
where $\mathcal{Q}_{k}$ is almost everywhere $\mathcal{C}^{1}$ over $\mathbb{R}^{2nd}$.
The Jacobian of mapping ${\bm{P}}_{1}({\bm{x}},{\bm{y}})-c\hat{\eta}_{k}u_{1,k}({\bm{x}},{\bm{y}})$ is
${\bm{J}}_{1c}({\bm{x}},{\bm{y}})=\begin{pmatrix}{\bm{W}}\otimes{\bm{I}}_{n}&0\\ 0&{\bm{I}}_{nd}\\ \end{pmatrix}-c\hat{\eta}_{k}\begin{pmatrix}0&\nabla^{2}\phi({\bm{y}})\\ 0&0\\ \end{pmatrix},$
where $\nabla^{2}\phi({\bm{y}})$ is well-defined over a full-measure set. And the Jacobian of the mapping ${\bm{P}}_{2}({\bm{x}},{\bm{y}})-c\hat{\eta}_{k}u_{2,k}({\bm{x}},{\bm{y}})$ is
${\bm{J}}_{2c}({\bm{x}},{\bm{y}})=\begin{pmatrix}{\bm{I}}_{nd}&0\\ 0&{\bm{W}}\otimes{\bm{I}}_{n}\\ \end{pmatrix}-c\hat{\eta}_{k}\begin{pmatrix}0&0\\ \tau\mathrm{Diag}(\nabla^{2}F_{1,\mathcal{B}_{1,k+1}}({\bm{x}}_{(1)}),\ldots,\nabla^{2}F_{d,\mathcal{B}_{d,k+1}}({\bm{x}}_{(d)}))&-\tau{\bm{W}}\otimes{\bm{I}}_{n}\\ \end{pmatrix},$
Henceforth, the Jacobian of ${\bm{P}}({\bm{x}},{\bm{y}})-c\mathcal{Q}_{k}({\bm{x}},{\bm{y}})$ is ${\bm{J}}_{2c}({\bm{x}},{\bm{y}}){\bm{J}}_{1c}({\bm{x}},{\bm{y}})$.
When ${\bm{W}}$ is invertible, then the determinant of ${\bm{J}}_{2}({\bm{x}},{\bm{y}}){\bm{J}}_{1}({\bm{x}},{\bm{y}})$ is a non-trivial polynomial of $c$ for almost everywhere $({\bm{x}},{\bm{y}})\in\mathbb{R}^{2nd}$. Leveraging similar proof techniques in Section A.1, we can derive that there exists a full-measure subset $\hat{\mathcal{S}}\subseteq\mathbb{R}_{+}$, such that for any $c\in\hat{\mathcal{S}}$, there exists a full-measure subset $\hat{\mathcal{K}}$ of $\mathbb{R}^{nd}\times\mathbb{R}^{nd}$ with the property that, whenever we choose $({\bm{x}}_{0},{\bm{y}}_{0})\in\hat{\mathcal{K}}$, it satisfies that $\{({\bm{x}}_{k},{\bm{y}}_{k})\}\cap\mathcal{Y}=\varnothing$. Consequently, the update scheme of vectorized sequence $({\bm{x}}_{k},{\bm{y}}_{k})$ can be reformulated as
[…]
Therefore, by applying Theorem 4.7 with $D_{f}=\partial f$, we can choose $\mathcal{S}=\hat{\mathcal{S}}\cap(0,\alpha_{c}),j=1,2$ and $\mathcal{K}=\hat{\mathcal{K}}\cap({\mathcal{X}}_{0}\times\mathbb{R}^{nd})$, such that the assertion in Theorem 4.8 holds.